# Two-stage Design for Failure Probability Estimation with Gaussian Process Surrogates

### v2

@@ Abstract
We tackle the problem of quantifying failure probabilities for expensive deterministic computer experiments with stochastic inputs. The computational cost of the computer simulation prohibits direct Monte Carlo (MC) and necessitates a surrogate model, turning the problem into a two-stage enterprise (surrogate training followed by probability estimation). Limited budgets create a design problem: how should expensive evaluations be allocated between and within the training and estimation stages? One may use the entire evaluation budget to sequentially train the surrogate through contour location (CL), with failure probabilities then estimated solely from the surrogate (we call it “surrogate MC”). But extended CL offers diminishing returns, and surrogate MC relies too stringently on surrogate accuracy. Alternatively, a partially trained surrogate may inform importance sampling, but this can provide erroneous results when budgets are limited. Instead we propose a two-stage design: starting with sequential CL, halting CL once learning has plateaued, then greedily allocating the remaining budget to MC samples with high classification entropy. Ultimately, we employ a “hybrid MC” estimator which leverages the trained surrogate in conjunction with the true responses observed in this second stage. Our unique two-stage design strikes an appropriate balance between exploring and exploiting, and outperforms alternatives, including both of the aforementioned approaches, on a variety of benchmark exercises. With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations, showcasing functionality with both shallow and deep Gaussian process surrogates, and deploying our method on a simulation of fluid flow around an airfoil.
Keywords: computer experiment, contour location, deep Gaussian process, emulator, entropy, importance sampling, Monte Carlo
@@ 1 Introduction
Quantifying the probability of failure in complex dynamic systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a deterministic black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$). The response is thresholded such that $f(\mathbf{x})>t$ indicates failure. The research objective is to estimate the failure probability,
@@
A rudimentary approach to solve Eq. (1) involves brute force Monte Carlo (MC): simply sample an exhaustively large set of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure. But this approach is prohibitively expensive if the computer simulation is costly (incurring long wall clock times), the probability of failure is rare (requiring enormous sample sizes), or both! In this work, we propose a new strategy for quantifying failure probabilities of complex computer simulations when the computation budget restricts evaluations of the simulator to only a few hundred, rendering traditional MC impractical.
In the face of expensive data, statistical models ($\hat{f}$, called “surrogates” or “emulators”) are essential. Surrogates are trained on a limited set of simulator evaluations in order to provide predictions with appropriate uncertainty quantification (UQ) at unobserved inputs. Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2018; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for nonstationary computer experiments (Sauer,, 2023). Notably, our contribution is agnostic to surrogate choice as long as effective UQ is provided, and is thus applicable to traditional GPs and Bayesian DGPs (Sauer et al.,, 2023b) and, potentially, Bayesian neural networks (MacKay,, 1995).
Choosing a surrogate model, like a GP or DGP, is only a first step. The utility of a surrogate boils down to two things: how it is trained and how it is used. For our objective (1), surrogate training should prioritize accurate classification of failures and surrogate deployment should focus on effective estimation of $\alpha$. The prevailing question is then: how can a limited budget of expensive simulator evaluations be optimally allocated between and within the training and estimation stages?
@@
To solve Eq. (1), surrogate accuracy in classifying failures is of primary importance. Not all training designs are created equally. Space-filling designs (Joseph,, 2016) may miss failure regions altogether, particularly when training budgets are limited. Instead, it is common to train a surrogate on a small initial design then pair it with a relevant acquisition function to select subsequent inputs for evaluation through the simulator. This iterative process of acquisition and surrogate updating can strategically hone in on the contour, all while minimizing evaluations of the expensive simulator. So-called “contour location” (CL) naturally outperforms space-filling counterparts. Acquisition functions seek to balance exploitation and exploration by leveraging both predicted distance from the contour and posterior uncertainty. Popular variations fall under expected improvement frameworks (Bichon et al.,, 2008; Ranjan et al.,, 2008), stepwise uncertainty reduction frameworks (Bect et al.,, 2012; Chevalier et al.,, 2014; Azzimonti et al.,, 2021), or both (Duhamel et al.,, 2023). Others utilize pass/fail classification entropy (Marques et al.,, 2018; Cole et al.,, 2023). Booth et al., (2024b) recently extended CL sequential designs to DGP surrogates. For small failure regions, contour location may be replaced with subset simulation (Au and Beck,, 2001) which initially targets a conservative threshold ($t^{\star}<t$) and iteratively refines it until the true threshold is found ($t^{\star}\rightarrow t$) (Dubourg et al.,, 2011; Huang et al.,, 2016; Hristov et al.,, 2019). Yet these approaches require Markov chain Monte Carlo (MCMC) sampling of cascading input distributions, which we intentionally avoid here.
@@
A well trained surrogate may be leveraged directly within an MC estimator, with surrogate predictions (which are of neglible cost) in place of true simulator evaluations. One could reasonably use the entire budget of simulator evaluations for contour location, then deploy the trained surrogate in such a “surrogate MC” estimator. Although this is common (e.g., Zhu and Du,, 2016; Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023), we find that surrogate learning of failure contours offers quickly diminishing returns. Furthermore, surrogate MC estimators place too much stock in the accuracy of the surrogate and funnel surrogate inaccuracies directly into the failure probability estimate.
@@
Another common approach splits the budget of simulator evaluations into two disjoint stages: some evaluations are used to train the surrogate and some are used for importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010). IS hinges on the construction of a bias distribution that intentionally focuses density in the failure region. Inputs are sampled from the bias distribution and are then utilized in a weighted average calculation that corrects for the disparity between the bias density and the original input density. Identifying an effective bias distribution in the face of limited training data is a hefty task (Tabandeh et al.,, 2022). Adaptive methods abound (e.g., Oh and Berger,, 1992; Kurtz and Song,, 2013; Dalbey and Swiler,, 2014; Bugallo et al.,, 2017; Cheng et al.,, 2023). Peherstorfer et al., (2016) propose what they termed “multifidelity importance sampling,” although we prefer the term “surrogate-informed importance sampling” (SIIS) to avoid confusion with multifidelity surrogate modeling (e.g., Park et al.,, 2017). In SIIS, a trained surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.3). SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024a; Renganathan,, 2024), and we view it as our primary competitor. Yet SIIS has several crucial drawbacks. We will delay a thorough discussion to Section 2.3, but the highlights are: (i) SIIS relies too heavily on an arbitrary bias distribution, (ii) it largely ignores the carefully trained surrogate, and (iii) it offers no guidance for allocating the budget between surrogate training and IS estimation.
@@
We prefer to avoid a bias distribution altogether, instead focusing on how we could optimally leverage a surrogate and an MC estimator in tandem. The answer lies in a two-stage design. First, we leverage sequential contour location to pin down the failure contour. At the start, it is essential to target inputs which provide the most insight into the entire failure contour (getting the most “bang for our buck” from early observations). As training progresses, successive acquisitions will offer less additional insight, indicated by apparent convergence of the surrogate MC estimator. Although we can not be sure of theoretical convergence with our limited budget, we propose a decision rule to recognize practical convergence (when surrogate learning from CL has saturated). We then spend the remaining simulation budget precisely on the Monte Carlo samples with the highest classification entropy, ultimately training the surrogate on observations from both stages. Our final failure probability estimate is a “hybrid MC” estimator which combines the observations from this second stage with surrogate predictions at unobserved samples.
@@
This manuscript is laid out as follows. Section 2, reviews GP/DGP surrogates, contour location, and importance sampling. Section 3.2 presents our two-stage design and failure probability estimator. Section 4 presents a variety of synthetic benchmarks. Finally, we deploy our method on a computer experiment of an RAE-2822 airfoil in Section 5 and conclude in Section 6.
@@ 2 Preliminaries
Denote single inputs as the row vector $\mathbf{x}$ of size $1\times d$ with response $y=f(\mathbf{x})$. Let $X_{n}$ represent the row-combined matrix of $n$-many inputs with corresponding response vector $\mathbf{y}_{n}$. Our objective is to quantify the failure probability from Eq. (1), namely $\mathrm{P}(f(\mathbf{x})>t)$ for $\mathbf{x}\sim p(\mathbf{x})$. Typical Monte Carlo with a budget of size $M$ provides the estimator
@@
This estimator is ineffective unless a sufficient number of failures are observed, necessitating very large $M$ particularly for small $\alpha$. With an expensive computer model, we simply cannot evaluate $f(\mathbf{x})$ thousands/millions/billions of times. Throughout, let $N$ denote the total budget of expensive simulator evaluations alotted. We will describe our sample efficient method for estimating failure probabilities in Section 3.2, targeting $N$ in the hundreds, after a quick review of preliminaries.
@@ 2.1 Gaussian process surrogates
A surrogate model $\hat{f}$, trained on a limited budget of simulations, may provide predictions at unobserved inputs in place of the true $f$. To enable our proposed methodology, the surrogate must also provide posterior predictive distributions. Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal distribution over the response. For any $n$ observed locations, our GP prior assumption is $\mathbf{y}_{n}\sim\mathcal{N}(\mu,\Sigma(X_{n}))$. The prior mean $\mu$ is often simplified to the zero vector after centering (which we will use moving forward), and the prior covariance $\Sigma(X_{n})$ is typically a function of inverse euclidean distances, e.g., $\Sigma(X_{n})^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$. See Santner et al., (2018); Rasmussen et al., (2006); Gramacy, (2020) for thorough reviews. Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singleton input $\mathbf{x}$ follow
$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x},\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(\mathbf{x},X_{n})^{\top},\end{array}$ (3)
@@
A natural alternative to true MC (2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
@@
Here we use the posterior mean of our GP in place of $f$. If the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
Figure 1: Left/Center: Herbie function with failure contour (red) and input distribution (white). Right: Entropy surface (yellow/high, purple/low) for a GP trained on a 100-point LHS (white circles).
Figure 2: Slice of Herbie function along $x_{2}=1.1$. GP was trained on 100-point LHS in 2 dimensions.
For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left and center panels of Figure 1 (we will revisit the right panel momentarily). We define a failure threshold of $t=1.065$ which demarcates four disjoint failure regions (red contours). We also define $p(\mathbf{x})$ using independent normal distributions centered at the origin (white contours). [^1] Full details of all simulated functions, thresholds, and distributions are provided in the Supplementary Material. We trained a GP surrogate using a random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n=100$ in two dimensions. Figure 2 shows the GP posterior prediction along the slice $x_{2}=1.1$. The posterior mean (dashed blue) and 95% confidence interval (blue shading) capture the true nonlinear surface (solid black) fairly well. Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures.” Thus any estimate reliant solely on this surrogate would underestimate the failure probability.

### v3

@@ Abstract
We tackle the problem of quantifying failure probabilities for expensive deterministic computer experiments with stochastic inputs under a fixed budget. The computational cost of the computer simulation prohibits direct Monte Carlo (MC) and necessitates a surrogate model, which may facilitate either a “surrogate MC” estimator or a surrogate-informed importance sampling estimator. We embrace the former, finding importance sampling too variable when budgets are limited, and propose a novel design strategy to effectively train a surrogate for the purpose of failure probability estimation. Existing works exhaust the entire evaluation budget on active learning through sequential contour location (CL), attempting to balance exploration with exploitation of the failure contour throughout the design, but we find exhaustive CL to be suboptimal. Instead we propose a novel two-stage surrogate design. In Stage 1, we conduct sequential CL to locate the failure contour. In Stage 2, once surrogate learning has saturated, we use a solely exploitative strategy – allocating the remaining evaluation budget to MC samples with the highest classification entropy to ensure they are classified correctly. We propose a stopping criterion to determine the transition between stages without any tuning. Our two-stage design outperforms alternatives, including exhaustive CL and surrogate-informed importance sampling, on a variety of benchmark exercises. With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations, showcasing functionality with both shallow and deep Gaussian process surrogates, and deploying our method on a simulation of fluid flow around an airfoil.
Keywords: active learning, contour location, deep Gaussian process, entropy, importance sampling
@@ 1 Introduction
Quantifying the probability of failure in complex systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a deterministic black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$). The response is thresholded such that $f(\mathbf{x})>t$ indicates failure. The research objective is to estimate the failure probability,
@@
from a small, fixed budget of simulator evaluations.
@@
A rudimentary approach to solve Eq. (1) involves brute force Monte Carlo (MC): sample a large number of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure. But this approach is prohibitively expensive if the computer simulation is costly and $\alpha$ is small, as in our target application. In the face of expensive data, statistical models ($\hat{f}$, called “surrogates” or “emulators”) are essential. Surrogates are trained on a limited set of simulator evaluations in order to provide predictions with appropriate uncertainty quantification (UQ) at unobserved inputs. Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2018; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for nonstationary computer experiments (Booth et al.,, 2024a). Notably, our contribution is agnostic to surrogate choice as long as effective UQ is provided, and is thus applicable to traditional GPs and Bayesian DGPs (Sauer et al.,, 2023b) and, potentially, Bayesian neural networks (MacKay,, 1995).
@@
The utility of a surrogate boils down to two things: how it is trained and how it is used. To solve Eq. (1), surrogate accuracy in classifying failures is of primary importance. Not all training designs are created equally. Space-filling designs (Joseph,, 2016) may miss failure regions altogether, particularly when training budgets are limited. Instead, it is common to train a surrogate on a small initial design then pair it with a relevant acquisition function to select subsequent inputs for evaluation through the simulator – a process called “active learning” (AL). The goal is to maximize learning from limited evaluations of the expensive simulator. Contour location (CL) is an AL variant which targets learning of a failure contour $\{\mathbf{x}\in\mathcal{X}\mid f(\mathbf{x})=t\}$. CL acquisition functions seek to balance exploitation and exploration by leveraging both predicted distance from the contour and posterior uncertainty. Popular variations fall under expected improvement frameworks (Bichon et al.,, 2008; Ranjan et al.,, 2008), stepwise uncertainty reduction frameworks (Bect et al.,, 2012; Chevalier et al.,, 2014; Azzimonti et al.,, 2021), or both (Duhamel et al.,, 2023). Others utilize pass/fail classification entropy (Marques et al.,, 2018; Cole et al.,, 2023). Booth et al., (2024c) recently proposed a CL sequential design for DGP surrogates. For small failure regions, contour location may be replaced with subset simulation (Au and Beck,, 2001) which initially targets a conservative threshold ($t^{\star}<t$) and iteratively refines it until the true threshold is found ($t^{\star}\rightarrow t$) (Dubourg et al.,, 2011; Huang et al.,, 2016; Hristov et al.,, 2019). Yet these approaches require Markov chain Monte Carlo (MCMC) sampling of cascading input distributions, which we intentionally avoid here.
@@
A well-trained surrogate may be leveraged directly within an MC estimator with surrogate predictions (which are of negligible cost) in place of true simulator evaluations, which we term “surrogate MC.” One could reasonably use the entire budget of simulator evaluations for contour location, then deploy the trained surrogate in such a “surrogate MC” estimator. Although this is common (e.g., Zhu and Du,, 2016; Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023), we find that surrogate learning of failure contours quickly plateaus, and that common surrogates fail to achieve sufficient classification accuracy under tightly restricted evaluation budgets.
@@
Another common approach splits the budget of simulator evaluations into two disjoint stages: some evaluations are used to train a surrogate and some are used for importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010). IS hinges on the construction of a bias distribution that intentionally focuses density in the failure region. Inputs are sampled from the bias distribution and are then utilized in a weighted average calculation that corrects for the disparity between the bias density and the original input density. Identifying an effective bias distribution in the face of limited training data is a hefty task (Tabandeh et al.,, 2022). Adaptive methods abound (e.g., Oh and Berger,, 1992; Kurtz and Song,, 2013; Dalbey and Swiler,, 2014; Bugallo et al.,, 2017; Cheng et al.,, 2023). Peherstorfer et al., (2016) propose what they termed “multifidelity importance sampling,” although we prefer the term “surrogate-informed importance sampling” (SIIS) to avoid confusion with multifidelity surrogate modeling (e.g., Park et al.,, 2017). In SIIS, a trained surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.3). SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024b; Renganathan,, 2024), and we view it as our primary competitor. Yet SIIS has several crucial drawbacks. We will delay a thorough discussion to Section 2.3, but the highlights are: (i) SIIS relies too heavily on an arbitrary bias distribution, (ii) it largely ignores the carefully trained surrogate, and (iii) it offers no guidance for allocating the budget between surrogate training and IS estimation.
Figure 1: Overview of proposed two-stage design for failure probability estimation. Stage 1 uses contour location to balance exploitation and exploration of $\{\mathbf{x}\in\mathcal{X}\mid f(\mathbf{x})=t\}$. Stage 2 solely exploits high uncertainty Monte Carlo samples from $p(\mathbf{x})$. For a fixed budget of $n+B$, our approach determines how to effectively allocate $n$ (for contour location) and $B$ (for Monte Carlo estimation).
@@
We prefer to avoid a bias distribution altogether, instead focusing on how we could optimally leverage a surrogate and an MC estimator in tandem. The answer lies in a two-stage surrogate design, represented in Figure 1 with details reserved for Section 3. First, in Stage 1 we conduct traditional contour location, balancing exploration and exploitation while locating the failure contour within domain $\mathcal{X}$. As training progresses, successive acquisitions will offer less additional insight, indicated by apparent convergence of the “surrogate MC” estimator. Although we can not be sure of theoretical convergence with our limited budget, we propose a stopping criterion to recognize practical convergence (when surrogate learning from CL has saturated). Once this stopping criterion is met, instead of continuing contour location or turning to importance sampling, we propose a novel “Stage 2” in which the remaining simulation budget is spent precisely on the Monte Carlo samples from $p(\mathbf{x})$ with the highest classification entropy. By observing the true $f(\mathbf{x})$ for samples where our surrogate is uncertain, we ensure these locations are classified correctly within our final MC estimate. Ultimately, surrogates trained on the observations acquired in these two unique stages provide more accurate and precise MC estimates of $\alpha$. While we prefer to view our contribution through the lens of surrogate design, our final failure probability estimate may be considered a “hybrid MC” estimator which combines the true simulator observations from Stage 2 with surrogate predictions at unobserved samples.
@@
This manuscript is laid out as follows. Section 2, reviews GP/DGP surrogates, contour location, and importance sampling. Section 3 presents our two-stage design and failure probability estimator. Section 4 presents a variety of synthetic benchmarks. Finally, we deploy our method on a computer experiment of an RAE-2822 airfoil in Section 5 and conclude in Section 6.
@@ 2 Preliminaries
Denote single inputs as the row vector $\mathbf{x}$ of size $1\times d$ with response $y=f(\mathbf{x})$, where $f$ represents the black-box computer simulation. Let $X_{n}$ represent the row-combined matrix of $n$-many inputs with corresponding response vector $\mathbf{y}_{n}$. Our objective is to quantify the failure probability from Eq. (1), namely $\mathrm{P}(f(\mathbf{x})>t)$ for $\mathbf{x}\sim p(\mathbf{x})$. Brute force Monte Carlo with a budget of size $M$ provides the estimator
@@
but this estimator is infeasible when $f$ is expensive. Let $N$ denote the total budget of expensive simulator evaluations allotted. We will describe our sample efficient method for estimating failure probabilities in Section 3, targeting $N$ in the hundreds, after a quick review of preliminaries.
@@ 2.1 Gaussian process surrogates
A surrogate model $\hat{f}$, trained on a limited budget of simulations, may provide predictions for unobserved inputs in place of the true $f$. To enable our proposed methodology, the surrogate must also provide uncertainty quantification through posterior predictive distributions. Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal distribution over the response. For any $n$ observed locations, our GP prior assumption is $\mathbf{y}_{n}\sim\mathcal{N}_{n}(\boldsymbol{\mu},\Sigma(X_{n}))$. The prior mean $\boldsymbol{\mu}$ is often simplified to the zero vector after centering (which we will use moving forward), and the prior covariance $\Sigma(X_{n})$ is typically a function of Euclidean distances, e.g., $\Sigma(X_{n})^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$. See Santner et al., (2018); Rasmussen et al., (2006); Gramacy, (2020) for thorough reviews. Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singleton input $\mathbf{x}$ follow
$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}_{1}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x},\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(\mathbf{x},X_{n})^{\top},\end{array}$ (3)
@@
A natural alternative to brute force MC (2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
@@
Here we use the posterior mean of our surrogate $\hat{f}$ in place of the true $f$. This estimator, however, inherits the bias of the surrogate – if the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
Figure 2: Left: Herbie function (yellow/high, purple/low) with true failure contour (red) and input distribution (white). Center/Right: Posterior predicted mean and entropy of a GP trained on an initial 20-point LHS (white circles) followed by 30 contour locating acquisitions (white triangles). Predicted contour in dashed red. Dashed line at $x_{2}=1.05$ indicates the slice shown in Figure 3.
Figure 3: Slice of Herbie function along $x_{2}=1.05$ (solid black) with failure threshold (dotted red). GP surrogate from Figure 2 shown in blue, with entropy in green along the $x$-axis. Yellow shaded regions indicate inputs which the surrogate would misclassify.
For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left panel of Figure 2 (we will revisit the other panels momentarily). We define a failure threshold of $t=1.065$ which demarcates four disjoint failure regions (red contours). We also define $p(\mathbf{x})$ using independent normal distributions centered at the origin (white contours). [^1] Full details of all simulated functions, thresholds, and distributions are provided in the Supplementary Material. We trained a GP surrogate using an initial random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n_{0}=20$ (white circles), followed by 30 contour locating acquisitions (white triangles, discussed in Section 2.2). Figure 3 shows the GP posterior along the slice $x_{2}=1.05$. The posterior mean (dashed blue) and 99% credible interval (blue shading) capture the true nonlinear surface (solid black) fairly well. Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures” or vice versa. While the predictions in these regions are reasonably accurate, their proximity to the contour makes precise classification difficult.